OpenStudy (anonymous):

Trigonometric Functions @ganeshie8

OpenStudy (anonymous):

\[Solve~the~following~equation~for~0\le~x~\le360\]

OpenStudy (anonymous):

\[\sin2x=1\]

OpenStudy (anonymous):

The first value is 45 but i'm not sure how to find the second value

OpenStudy (freckles):

let u=2x then x=u/2 so you have 0<=u/2<=360 solving for u we get 0<=u<=720 so first step solve sin(u)=1 on the interval [0,720]

OpenStudy (freckles):

then once you find u then you can replace u with 2x and solve for x

OpenStudy (anonymous):

\[u=90\]\[x=\frac{ 90 }{ 2 }\]\[x=45\]is it correct? @freckles

OpenStudy (freckles):

ok but there are other u in the interval such that sin(u)=1

OpenStudy (freckles):

you have sin(u)=1 when u=90,90+360 and so 2x=90,90+360 so x=90/2 or x=(90+360)/2

OpenStudy (freckles):

@MARC_

OpenStudy (anonymous):

\[x=45,225\]

OpenStudy (anonymous):

Thank you @freckles :)

OpenStudy (freckles):

np